Winter-2018 Phys 2701-01 Name: 20) S 10m/s, road you raise or lower only 14 m in elevation (vertical) uppose your car turn to Carlos Bee drive with the initial velocity of, Vi- the road has a 14% grade, meaning that for every 100 m horizontal travel a) What will be your acceleration down the hill of the road, if you did not press on accelerator of your car? b) What will be your final velocity at the bottom of the hill with the road distance of 300m and initial velocity of 10m/s? (Assume, no accelerator is pressed during down the hill travell). c) How long the car took to reach the bottom of the hill, with V,-10m/s and distance of 300m?
Answer #1

20. tan(theta) = 0.14

theta = 7.97 deg


(A) along the incline, Fnet = m g sin(theta) ....down the incline

a = F / m = g sin7.97

a = 9.8 x sin7.97 = 1.36 m/s^2 (down the incline)

(B) Applying vf^2 - vi^2 = 2 a d

v^2 - 10^2 = 2(1.36)(300)

v = 30.25 m/s


(c) vf = vi + a t

30.25 = 10 + 1.36 t

t = 14.9 sec